# Testing the dark side of neutrino oscillations with the solar neutrino fog at Dark Matter experiments

### v1

@@ V.1 LMA-Dark in muon and tau neutrino sector
Turning now to our results. We show in fig. 2 the constraints derived from the XENONnT and PandaX-4T data sets Aprile et al. (2024a); Bo et al. (2024) in the plane for a mediator with equal couplings to muon and tau neutrinos and equal couplings to up and down quarks compared to the LMA-Dark region from Esteban et al. (2018). The data from current experiments only excludes the LMA-Dark region above mediator masses of MeV.
As current data does not fully exclude LMA-Dark in the muon and tau sector, we consider possible future DM direct detection experiments and test their sensitivity. PandaX-4T Abdukerim et al. (2025) is proposed to be upgraded to the PandaX-xT experiment, and will feature a 43 t liquid xenon target. The experiment is expected to conservatively have an efficiency as good as the current generation PandaX-4T Abdukerim et al. (2025), with the possibility of improved efficiency and lower backgrounds. XLZD is the planned upgraded detector of the XENON and LZ experiments Aalbers et al. (2016, 2024). In the conservative plan, XLZD aims to have a 60 t liquid xenon target. Both future xenon experiments expect to reach a 200 t-yr exposure by the end of their lifetime. Darkside-20k Aalseth et al. (2018) and ARGO arg are planned DM experiment which would feature a 48 t and a 300 t liquid argon target and they are expected to achieve a threshold of 0.5 keVnr, when utilizing the ionization only channel Atzori Corona (2025). The number of CEvNS events in future xenon and argon detectors is similar, due to the similar nuclear masses, we therefore focus on the xenon experiments in the following but similar results also apply to future argon experiments.
@@
If the experimental parameters improve in future iterations of DM direct detection experiments LMA-Dark can be ruled out with lower exposure. In fig. 3 we show the required values of and as a function of exposure which excludes LMA-D at (see appendix A for the results at ). Depending on the background level and the systematic uncertainty on the signal and background events, the required exposure can be reduced to . Lowering the background requires to reduce the radioactivity on the cathode electrode which dominates the number of background events for the US2 data of PandaX-4T Bo et al. (2024) or to reduce the accidental coincidence events for XENONnT Aprile et al. (2024a).
Given that the current signal to background ratio of the total number of events is 0.3 for PandaX-4T with varying signal-to-background ratios across the bins between 0.25 to 0.46, decreasing the systematic uncertainty on the signal is less impactful than decreasing the uncertainty on the background . The systematic uncertainty of the signal comes from the experimental data selection and the 8B solar neutrino flux. The latter uncertainty could potentially be reduced in the future as upcoming experiments, in particular solar neutrino measurements at DUNE Capozzi et al. (2019); Abi et al. (2020); Meighen-Berger et al. (2025), can measure the 8B flux with an improved precision of 1.6% Meighen-Berger et al. (2025) compared to the current uncertainty of 5% Aharmim et al. (2013); Abe et al. (2024); Agostini et al. (2018).
If the assumption of equal couplings to up-and down quarks is softened, the excluded regions for a mediator of mass MeV coupling with equal couplings to to muon and tau neutrinos is shown fig. 4. A future xenon experiment with 200 t-yr exposure will narrow the allowed regions drastically. The slope of the bands in the is given by . Therefore, materials with a different neutron to proton ratios are required to break the degeneracies in the parameter space and shrink the allowed regions. Note that oscillation data using neutrino propagation through the Earth and the Sun only slightly narrows down this parameter space Denton and Gehrlein (2022) as the neutron to proton ratios differ only by a factor of between 1.05 for the Earth and 0.3 in the Sun for the peak of the 8B flux Bahcall et al. (2005) compared to of Xenon. We show in fig. 4 the LMA and LMA-Dark degenerate regions in the Sun and in the Earth. Note that there is a degenerate point which cannot be distinguished from the case of no new physics in oscillation experiments.
@@ V.2 Distinguishing new physics scenarios in the muon and tau neutrino sectors
To illustrate the power of DM experiments to disentangle a new light mediator coupling to muon and tau neutrinos or just one flavor, we assume that LMA-Dark is realized in nature with a light mediator and a coupling corresponding to the LMA-Dark region with . We test then a scenario where the light mediator couples only to muon or only to tau neutrinos with the coupling strength corresponding to the LMA-Dark region. The distinction power between these two scenarios parametrized as is shown in fig. 5. We see that current data can only distinguish between these scenarios for mediator masses above 15 MeV at . The results for new interactions of muon or of tau neutrinos only are very similar due to the similar fluxes of these flavors. For clarity, we show results for future experimental configurations for the couplings to muon neutrinos only. An exposure of 50 t-yr for a PandaX-4T like experiment is required to distinguish between these scenarios at down to mediator masses excluded by cosmology. Around a mediator mass of MeV, the number of events in our scenarios is close to the SM value, making it challenging to exclude new physics and leading to a reduced sensitivity.
@@ V.3 LMA-Dark in the electron neutrino sector
If LMA-Dark is realized in the electron neutrino sector with couplings to quarks, i. e. and all other NSI parameters are zero, DM experiments can probe this region as well. Recently, the first observations of CEvNS with reactor neutrinos has been announced from the Dresden-II experiment Colaresi et al. (2022, 2021) and CONUS+ Ackermann et al. (2025) which were analyzed in Coloma et al. (2022b); Aristizabal Sierra et al. (2022); Liao et al. (2022); Chattaraj et al. (2025a); De Romeri et al. (2025); Atzori Corona et al. (2025). We show their constraints in fig. 6 together with the current constraints from PandaX-4T and XENONnT. While current XENONnT and PandaX-4T data only probes the LMA-Dark region down to mediator masses of 10 MeV, similar as for LMA-Dark in the muon and tau sector due to the similar neutrino flux, data from the Dresden-II experiment fully excluded the region, independent of the choice of quenching factor and the CONUS+ results cover the vast majority of parameter space as well. Future PandaX-4T like experiments can provide complementary probes to reactor CEvNS experiments but requires an exposure of at least 600 t-yr to exclude LMA-Dark at .
@@ VI Conclusions
The first observation of the neutrino background at DM direct detection experiments allows to probe new contributions to the neutrino scattering cross section at low energies. Of particular interest is a region of parameter space of new neutrino interactions which leads to the LMA-Dark degeneracy in neutrino oscillations. This degeneracy of the oscillation Hamiltonian corresponds to a change in the mass ordering in the presence of new vectorial neutrino interactions with strength of order of the Fermi constant. Here we have shown that DM direct detection experiments have an unparalleled possibility to probe a yet unexplored region of LMA-Dark parameter space, namely equal couplings of a new light mediator to muon and tau neutrinos. This is due to the presence of all three neutrino flavors in the 8B solar neutrino flux and their subsequent detection in the low-energy CEvNS scattering process.
We have shown that current solar neutrino CEvNS data from XENONnT and PandaX-4T does not yet fully exclude the LMA-Dark parameter space for equal couplings to muon and tau neutrinos, however, assuming similar experimental details, xenon based DM experiments with an exposure of 125 t-yr are sufficient to achieve this goal at 3. Finally, we have also shown that DM experiments provide a complementary way to probe LMA-Dark in the electron sector, that is a light mediator coupling to quarks and electron neutrinos, compared to reactor neutrino CEvNS experiments.
The remaining LMA-Dark regions of parameter space are equal interactions of muon and tau neutrinos with electrons mediated by a low mass vector boson can be probed in the future in large scale solar neutrino-electron scattering experiments like JUNO, DUNE, or HyperKamiokande or DM direct detection experiments.

### v2

@@ V.1 LMA-Dark in muon and tau neutrino sector
We turn now to our results. We show in fig. 2 the constraints derived from the XENONnT and PandaX-4T data sets Aprile et al. (2024a); Bo et al. (2024) in the plane for a mediator with equal couplings to muon and tau neutrinos and equal couplings to up and down quarks compared to the LMA-Dark region from Esteban et al. (2018). The data from current experiments only excludes the LMA-Dark region above mediator masses of MeV.
As current data does not fully exclude LMA-Dark in the muon and tau sector, we consider possible future DM direct detection experiments and test their sensitivity. PandaX-4T Abdukerim et al. (2025) is proposed to be upgraded to the PandaX-xT experiment, and will feature a 43 t liquid xenon target. The experiment is expected to conservatively have an efficiency as good as the current generation PandaX-4T Abdukerim et al. (2025), with the possibility of improved efficiency and lower backgrounds. XLZD is the planned upgraded detector of the XENON and LZ experiments Aalbers et al. (2016, 2024). In the conservative plan, XLZD aims to have a 60 t liquid xenon target. Both future xenon experiments expect to reach a 200 t-yr exposure by the end of their lifetimes. Darkside-20k Aalseth et al. (2018) and ARGO arg are planned DM experiments which would feature a 48 t and a 300 t liquid argon target and they are expected to achieve a threshold of 0.5 keVnr, when utilizing the ionization only channel Atzori Corona (2025). The number of CEvNS events in future xenon and argon detectors is similar, due to the similar nuclear masses, and we therefore focus on the xenon experiments in the following but similar results also apply to future argon experiments.
@@
If the experimental parameters improve in future iterations of DM direct detection experiments LMA-Dark can be ruled out with lower exposure. In fig. 3 we show the required values of and as functions of exposure which excludes LMA-D at (see appendix A for the results at ). Depending on the background level and the systematic uncertainty on the signal and background events, the required exposure can be reduced to . Lowering the background requires reducing the radioactivity on the cathode electrode which dominates the number of background events for the US2 data of PandaX-4T Bo et al. (2024) or to reduce the accidental coincidence events for XENONnT Aprile et al. (2024a).
Given that the current signal-to-background ratio of the total number of events is 0.3 for PandaX-4T with varying signal-to-background ratios across the bins between 0.25 to 0.46, decreasing the systematic uncertainty on the signal is less impactful than decreasing the uncertainty on the background . The systematic uncertainty of the signal comes from the experimental data selection and the 8B solar neutrino flux. The latter uncertainty could potentially be reduced in the future as upcoming experiments, in particular solar neutrino measurements at DUNE Capozzi et al. (2019); Abi et al. (2020); Meighen-Berger et al. (2025), can measure the 8B flux with an improved precision of 1.6% Meighen-Berger et al. (2025) compared to the current uncertainty of 5% Aharmim et al. (2013); Abe et al. (2024); Agostini et al. (2018).
If the assumption of equal couplings to up and down quarks is softened, the excluded regions for a mediator of mass MeV coupling with equal couplings to muon and tau neutrinos is shown fig. 4. A future xenon experiment with 200 t-yr exposure will narrow the allowed regions drastically. The slope of the bands in the plane is given by . Therefore, materials with a different neutron to proton ratios are required to break the degeneracies in the parameter space and shrink the allowed regions. Note that oscillation data using neutrino propagation through the Earth and the Sun only slightly narrow down this parameter space Denton and Gehrlein (2022) as the neutron to proton ratios differ only by a factor of between 1.05 for the Earth and 0.3 in the Sun for the peak of the 8B flux Bahcall et al. (2005) compared to of Xenon. We show in fig. 4 the LMA and LMA-Dark degenerate regions in the Sun and in the Earth. Note that there is a degenerate point which cannot be distinguished from the case of no new physics in oscillation experiments.
@@ V.2 Distinguishing new physics scenarios in the muon and tau neutrino sectors
To illustrate the power of DM experiments to disentangle a new light mediator coupling to muon and tau neutrinos or just one flavor, we assume that LMA-Dark is realized in nature with a light mediator and a coupling corresponding to the LMA-Dark region with . We test then a scenario where the light mediator couples to only muon or only tau neutrinos with the coupling strength corresponding to the LMA-Dark region. Since we assumed and the transition probabilities and are identical leading to the same results for a mediator coupling to either only muon neutrinos or only tau neutrinos. The distinction power between these two scenarios parametrized as is shown in fig. 5. We see that current data can only distinguish between these scenarios for mediator masses above 15 MeV at . An exposure of 50 t-yr for a PandaX-4T-like experiment is required to distinguish between these scenarios at down to mediator masses excluded by cosmology. Around a mediator mass of MeV, the number of events in our scenarios is close to the SM value, making it challenging to exclude new physics and leading to a reduced sensitivity.
@@ V.3 LMA-Dark in the electron neutrino sector
If LMA-Dark is realized in the electron neutrino sector with couplings to quarks, i. e. and all other NSI parameters are zero, DM experiments can probe this region as well. Recently, the first observations of CEvNS with reactor neutrinos has been announced from the Dresden-II experiment Colaresi et al. (2022, 2021) and CONUS+ Ackermann et al. (2025) which were analyzed in Coloma et al. (2022b); Aristizabal Sierra et al. (2022); Liao et al. (2022); Chattaraj et al. (2025a); De Romeri et al. (2025); Atzori Corona et al. (2025). We show their constraints in fig. 6 together with the current constraints from PandaX-4T and XENONnT. While current XENONnT and PandaX-4T data only probes the LMA-Dark region down to mediator masses of 10 MeV, similar to LMA-Dark in the muon and tau sector due to the similar neutrino flux, data from the Dresden-II experiment fully excluded the region, independent of the choice of quenching factor and the CONUS+ results cover the vast majority of parameter space as well. Future PandaX-4T-like experiments can provide complementary probes to reactor CEvNS experiments but require an exposure of at least 600 t-yr to exclude LMA-Dark at .
@@ VI Conclusions
The first observation of the neutrino background at DM direct detection experiments allows to probe new contributions to the neutrino scattering cross section at low energies. Of particular interest is a region of parameter space of new neutrino interactions which leads to the LMA-Dark degeneracy in neutrino oscillations. This degeneracy of the oscillation Hamiltonian corresponds to a change in the mass ordering in the presence of new vectorial neutrino interactions with strength of order the Fermi constant. Here we have shown that DM direct detection experiments have an unparalleled possibility to probe a yet unexplored region of LMA-Dark parameter space, namely equal couplings of a new light mediator to muon and tau neutrinos. This is due to the presence of all three neutrino flavors in the 8B solar neutrino flux coming from the transition probabilities of electron neutrinos to muon and tau neutrinos in addition to the non-vanishing electron neutrino survival probability Wolfenstein (1978) and their subsequent detection in the low-energy CEvNS scattering process.
We have shown that current solar neutrino CEvNS data from XENONnT and PandaX-4T does not yet fully exclude the LMA-Dark parameter space for equal couplings to muon and tau neutrinos, however, assuming similar experimental details, xenon-based DM experiments with an exposure of 125 t-yr are sufficient to achieve this goal at 3. Finally, we have also shown that DM experiments provide a complementary way to probe LMA-Dark in the electron sector, that is a light mediator coupling to quarks and electron neutrinos, compared to reactor neutrino CEvNS experiments.
The remaining LMA-Dark regions of parameter space are equal interactions of muon and tau neutrinos with electrons mediated by a low mass vector boson and can be probed in the future in large-scale solar neutrino-electron scattering experiments like JUNO, DUNE, or HyperKamiokande or DM direct detection experiments.